OpenStudy (anonymous):

Sarah went to an electronics store, purchased 3 CDs and 5 DVDs, and spent $137. Michelle went to the same store, bought 2 CDs and 7 DVDs, and spent $161. How much does each CD sell for at the store? Answer $19 $14 $18 $13

OpenStudy (anonymous):

3x+5y=137 2x+7y=161 x..CD and y..DVD

OpenStudy (anonymous):

21x+35y=959 10x+35y=805 ------------- 11x=154 x=14